OpenStudy (anonymous):

Find the general term an for the geometric sequence a1= -4 and a2 = 20

OpenStudy (anonymous):

\[a _{n} a _{1} = -4 a _{2} = 20 \]

OpenStudy (anonymous):

@satellite73 help?

OpenStudy (amistre64):

\[a_1~r = a_2\]what is r?

OpenStudy (anonymous):

he didn't give me R he only gave me a1 and a2

OpenStudy (amistre64):

.... then you might want to plug that a1 and a2 into the formula and SOLVE for r

OpenStudy (anonymous):

what is that formula?

OpenStudy (anonymous):

the a1+ (n-1)d

OpenStudy (amistre64):

no, we do not want to find an arithmetic sequence with a common difference (d) we want to find an geometric sequence with a ratio (r) a geometric sequence is the result of multiplying the same value (r) to each new an in order to determine the next value. therefore, you are given 2 values, a1 and a2. By applying the definition of a geometric sequence, the value of a2 must be a1 times r

OpenStudy (amistre64):

therefore: \[a_1~r=a_2\] what is the common ratio we want to determine? what is the value of "r" ?

OpenStudy (anonymous):

well to get a2 to be 20, you can multiply -4 by -5

OpenStudy (amistre64):

correct :) so we have 2 ways we can express the \(a_n\) term. recurrsively, and also explicitly which way do you need to express it?

OpenStudy (anonymous):

he doesn't state in which way to express it

OpenStudy (amistre64):

recurrsively is defining the n+1 in terms of n\[a_{n+1}=a_{n}~r~;~a_1 = k\] the other way ...

OpenStudy (amistre64):

explicitly stated:\[a_n=a_1 ~r^{n-1}\]

OpenStudy (anonymous):

Let me look and see how they are showing this in the chapter, it's only suppose to be an algebra class

OpenStudy (amistre64):

ok, im assuming they want the explicit then :)

OpenStudy (anonymous):

okay, your way better at this then me. do we assume n = 1 ?

OpenStudy (amistre64):

n is the value of the nth term of the sequence that we would want to fine, its best to leave it as the variable in the equation so that we can fill it in as needed

OpenStudy (anonymous):

okay leaving the equation to look like a n = -4(-5) n-1

OpenStudy (amistre64):

spose we wanted to know the 156th term of the sequence the recurrsion equation would take some time to get all the way up to 156 the explicit equation allows us a more convient method:\[a_{156}=a_1~r^{156-1}\]

OpenStudy (amistre64):

yes:\[a_n=-4~(-5)^{n-1}\]

OpenStudy (anonymous):

so for the general term is an= -4(-5) n-1

OpenStudy (amistre64):

yes, make sure you exponent that (n-1) tho.

OpenStudy (anonymous):

have time for one more?

OpenStudy (amistre64):

maybe, hard to tell the future :)

OpenStudy (anonymous):

Write the terms of their series and find their sums \[\sum_{k=1}^{4} (k+5)^{2} \]

OpenStudy (amistre64):

well, id write out (k+5)^2 .... enough times to calculate k=1,2,3,4 ; so four times

OpenStudy (amistre64):

then for each time youve written it, fill in a new k value and evaluate the setup

OpenStudy (amistre64):

1: (k+5)^2 2: (k+5)^2 3: (k+5)^2 4: (k+5)^2 1: (1+5)^2 2: (2+5)^2 3: (3+5)^2 4: (4+5)^2 1: (6)^2 2: (7)^2 3: (8)^2 4: (9)^2

OpenStudy (anonymous):

on our homework we had the same equation but the top was 4 and k=3, he had us figuring it out by k= 3^2-3= 9-3 = 6 4^2 -4 = 16-4 = 12 12+6 = 18

OpenStudy (amistre64):

ok, then the question is asking you to find partial sums, or a running cumulation of the sums of the sequnce that is generated

OpenStudy (anonymous):

I had to find the series and it's sum of the equation.

OpenStudy (amistre64):

a1: (6)^2 a2: (7)^2 a3: (8)^2 a4: (9)^2 S1 = a1 S2 = a1 + a2 S3 = a1 + a2 + a3 S4 = a1 + a2 + a3 + a4

OpenStudy (anonymous):

so the series would be 36,49, 64, and 81

OpenStudy (amistre64):

the "sequence" is 36,49,64, and 81 the "series" is: 36 + 49 + 64 + 81

OpenStudy (anonymous):

so 230

OpenStudy (amistre64):

sounds about right

OpenStudy (anonymous):

okay so 4 ∑k=1(k+5)2 how do we solve for that?

OpenStudy (amistre64):

you just did, that is what we just went thru

OpenStudy (amistre64):

4 ∑k=1(k+5)2 means: (1+5)^2 + (2+5)^2 + (3+5)^2 + (4+5)^2 its just more concise to write it as: 4 ∑k=1(k+5)2 to express the same thing

OpenStudy (anonymous):

ooh okay, figured that out now. the way he set up the question threw me off. Okay so the terms of the sequence is 36,49,64,81

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

okay, I am sorry I am not trying to make this harder on you, I am just so lost in this class.

OpenStudy (amistre64):

if its new material, then there is a bit of a learning curve :) im just used to the notations better

OpenStudy (anonymous):

high school algebra sure has changed since I went. I never had to do things like this so it is a HUGE learning curve.

OpenStudy (amistre64):

i was 20 years out of high school and went to college to get a masters in math. my first class i found out that i had retained half of my school math, forgot about 1/4 of it; and there was 1/4 i had never even come across

OpenStudy (anonymous):

it's been about 5 year since high school, getting my degree in the medical field. Thank you so much for your help and taking the time . I have two more problem to see if anyone can help with =)

OpenStudy (amistre64):

good luck :) ill be around ....

OpenStudy (anonymous):

thank you.